​(DuPont analysis)  Dearborn Supplies has total sales of

$ 191 ​million, assets of $ 90 ​million, a return on equity of 31 ​percent, and a net profit margin of 7.9 percent.

What is the​ firm's debt​ ratio?
Answer #1
Total asset turnover = Sales/total assets
Total asset turnover = 90/191
Total asset turnover = 0.47
DUPONT
ROE = Net profit margin*total asset turnover*equity multiplier
0.31=0.079*0.47*Equity multiplier

Equity multiplier= 8.349

A/E = 8.349

E/A = 1/8.349=0.11977

D/A = 1-E/A =1-0.11977=0.88